2 g of He, 3 g of N; and 4 g of Ar were introduced into a 15 dm
1
Expert's answer
2021-09-09T21:38:01-0400

Moles of He "=\frac{2}{4.00}=0.5moles"

Moles of N"=\frac{3}{14.00}=0.214"


Moles Ar"=\frac{4}{39.948}=0.102moles"


Average moles "=\frac{0.5+0214+0.102}{3}=0.272moles"



Molarity "=\frac{0.272}{15}=0.018M"